OpenStudy (anonymous):

Help please, best response will be awarded. Mickey deposited $782.11 in a savings account that earns 4.3% simple interest. What is Mickey’s account balance after 16 years? $815.74 $1,320.20 $538.09 $5,380.92

OpenStudy (phi):

what is the formula for simple interest ?

OpenStudy (nurali):

Simple Interest = (PTR)/100 Simple Interest = (782.11*4.3*16)/100 Simple Interest = 538.09 And AMOUNT =PRINCIPLE + SIMPLE INTEREST =538.09 + 782.1 =1,320.20